[Brahms-l] Shift report 20050127 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Jan 27 2005 - 06:22:05 EST
1. RHIC delivered 3 hours of beam.  They had some problems between stores, so it took them a while to get going again.  When they came back up the rates were high.  

2. I noticed that we collected 2Mil trig 3 events for the MRS 90deg B2255 data, so I ramped D5 to 700amps.  I collected around 170k trig 3 events, but I had problems stopping the run.  So these event are probably not cataloged correctly.  I am not sure what the status of this data is.  The run numbers are 12534 and 12529.  

3. There are a number of HV channels that have the default values different than the target values.  I have been working under the impression that the target values are the correct ones.  

4. DAQ DID NOT STOP PROPERLY FOR THE LAST TWO RUNS I TOOK!!!!

5. A channel in the TOF2 still looks "hot".

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 27 06:22:24 2005

This archive was generated by hypermail 2.1.8 : Thu Jan 27 2005 - 06:22:43 EST